OpenStudy (h0pe):

An eccentric baseball card collector wants to distribute her collection among her descendants. If she divided her cards among her 17 great-great-grandchildren, there would be three cards left over. If she divided her cards among her 16 great-grandchildren, there would be 10 cards left over. If she divided her cards among her 11 grandchildren, there would be 4 cards left over. If she divided her cards among her 7 children, there would be no cards left over. What is the smallest possible number of cards in her collection?

OpenStudy (rational):

tried anything yet ?

OpenStudy (h0pe):

No, not really, but I'm guessing I need to solve this with mods?

OpenStudy (rational):

Let \(x\) = number of cards set up four congruences and solve \(x\) \[x\equiv 3\pmod{17}\\~\\x\equiv 10\pmod{16}\\~\\x\equiv 4\pmod{11}\\~\\x\equiv 0\pmod{7}\]

OpenStudy (h0pe):

Should I substitute? like \[x=17x+3=16x+10\] or something?

OpenStudy (h0pe):

and \[x=11x+4=7x\]Well they have to have different variables right?

OpenStudy (rational):

also notice starting from bottom makes it easy, start with x = 7k you may use congruence properties instead

OpenStudy (h0pe):

so \[x=7k\]\[x=11a+4\]\[x=16b+10\]\[x=17c+3\]?

OpenStudy (rational):

looks good, but i prefer working with congruences as they are simpler

OpenStudy (rational):

anyways, as u can see u can work it exact same way as we did in the previous problem simply plugin x=7k in second equation and solve k

OpenStudy (h0pe):

so \[x=7k=11a+4\]? How would you do it with congruences?

OpenStudy (rational):

\[7k\equiv 4\pmod{11}\] heard of "inverses" ?

OpenStudy (rational):

what is the inverse of 7 in mod 11 ?

OpenStudy (h0pe):

it's 8

OpenStudy (anonymous):

google chinese remainder theorem and you will get a method for solving this

OpenStudy (is3535):

8 is your answer medal me plzz

OpenStudy (h0pe):

it's not..

OpenStudy (rational):

\[7k\equiv 4\pmod{11}\] multiply the inverse of 7 both sides and solve \(k\)

OpenStudy (h0pe):

like k\[k≡32 (\mod11)\]?

OpenStudy (rational):

Yes, which is same as \(k \equiv 10 \pmod{11}\)

OpenStudy (rational):

same as \(k = 11t + 10\) so \(x = 7k = 7(11t+10)=77t+70\)

OpenStudy (rational):

plug that in next congruence and solve \(t\)

OpenStudy (h0pe):

What do you mean by the next congruence?

OpenStudy (rational):

http://gyazo.com/9a268d4a77b190da48ffd1748ca13f10

OpenStudy (rational):

from the first two congruences we have \[x=77t+70\] plug this in third congruence and solve \(t\)

OpenStudy (h0pe):

The next congruence is \[x=16b+10\]

OpenStudy (rational):

\(x = 16b+10\) is same as \[x\equiv 10\pmod{16}\]

OpenStudy (rational):

simply plugin \(x=77t+70\) above and reduce

OpenStudy (h0pe):

so \[77t+70=10(\mod 16)\]

OpenStudy (h0pe):

do i multiply the inverse of 10 mod 16 on both sides?

OpenStudy (rational):

first reduce

OpenStudy (rational):

\(77t+70\equiv 10 \pmod{16}\) \(77t \equiv -60\pmod{16}\) \(77t \equiv 4\pmod{16}\) yes ?

OpenStudy (h0pe):

oh right

OpenStudy (h0pe):

so then now the inverse?

OpenStudy (rational):

also 77 is same as -3 in mod 16, so \(77t \equiv 4\pmod{16}\) reduces to \(-3t \equiv 4\pmod{16}\)

OpenStudy (rational):

now what is the inverse of -3 in mod 16

OpenStudy (rational):

in other words, what number you need to multiply by "-3" to get 1 ?

OpenStudy (h0pe):

-11

OpenStudy (h0pe):

so that gives me \[t≡-44(\mod16)\]?

OpenStudy (rational):

Notice that -11 is same as 5

OpenStudy (rational):

-44 is correct, reduce it and make it look simpler

OpenStudy (h0pe):

that's basically \[t≡4(\mod16)\] isn't it

OpenStudy (h0pe):

which is \[t=16x+4\]

OpenStudy (h0pe):

is that correct?

OpenStudy (rational):

\(t = 16m+4\)

OpenStudy (rational):

so \(x = 77t+70 = 77(16m+4)+70 = 1232m + 378\) plug that in next congruence and solve \(m\)

OpenStudy (h0pe):

So \[1232m=-375(\mod17)\]and then\[m=-5625(\mod17)\]

OpenStudy (h0pe):

Which is basically \[m=2(\mod17)\]

OpenStudy (rational):

Yes, which is same as \(m = 17q + 2\)

OpenStudy (rational):

so \(x = 1232m+378 = 1232(17q+2)+378 = 20944q +2842 \)

OpenStudy (rational):

so the smallest possible collection of cards is \(2842\)

OpenStudy (h0pe):

Thanks!

OpenStudy (rational):

Also notice \[20944 = 7*11*16*17\]

OpenStudy (rational):

the whole thing we did just now can be done in a single line using chinese remainder theorem

OpenStudy (h0pe):

woah :o better google that

OpenStudy (rational):

i suggest you wait till your professor teaches you that

OpenStudy (h0pe):

alright, thank you!

OpenStudy (rational):

np